OpenStudy (anonymous):

the sum of two numbers is 11. If one number is subtracted from the other, the result is -1. Find the numbers

OpenStudy (anonymous):

a + b = a - b = (b-a will give same result with a,b reversed)

OpenStudy (anonymous):

6,5.